# Generating $k$ EPR-pairs from an $n$-party resource state

### v4

@@ 1.1 Generating EPR-pairs from a resource state
Quantum communication networks combine several quantum computers to enable them to solve interesting tasks from cryptography, communication complexity, distributed computing etc. Building a large-scale quantum communication network is a daunting task that will take many years, but networks with a few small quantum computers are under construction and may start to appear in the next few years Wehner et al. (2018).
@@
These networks are either based on channels that physically communicate quantum states, or rely on classical communication in tandem with shared entanglement, or a combination of both. Communication over classical channels cannot increase entanglement, so in the absence of quantum channels we have to rely on prior entangled states. For example, if two parties share an EPR-pair, $\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$, then one party can transmit (“teleport”) a qubit to the other via two classical bits of communication, consuming the EPR-pair in the process Bennett et al. (1993). If we want to enable many qubits to be sent in this way, then we could start with an entangled state where each pair among the $n$ parties shares its own EPR-pair. This would allow any pair to exchange a qubit, but would require us to start with a rather large initial entangled state of $\binom{n}{2}$ EPR-pairs, and each of the $n$ parties would need to hold $n-1$ qubits (see Figure 1 for $n=4$).
@@ 1.2 Motivation
Multi-party resource states are foundational for tasks such as distributed quantum computing Raussendorf et al. (2003); Raussendorf and Briegel (2001); Meter (2014), quantum secret sharing Hillery et al. (1999); Meter (2014), and multi-party quantum protocols Cleve and Buhrman (1997); Fitzi et al. (2001). One of the most utilized resources in quantum networking is EPR-pairs shared between two parties. Our model aims to capture scenarios involving multiple parties, where it is not known in advance between which parties the quantum resources should be deployed. This situation is similar to a call center, which must dynamically manage incoming calls. Our study concentrates on developing resource states that provide the flexibility to determine the call/connection structure after the resource state has been created.
@@
In this theoretical work we use several simplifying assumptions, most notably that EPR-pairs suffice as the primary resource required for applications. The interesting follow-up work Claudet et al. (2023); Cautés et al. (2024), discussed at the end of this section, subsequently removed this assumption.
@@ 1.3 Our results 1: constructions of $k$-pairable resource states
Our protocols for generating EPR-pairs can be made fault-tolerant if each party encodes their $m$ qubits by some stabilizer-type quantum code Gottesman (1998b). Importantly, the encoded versions of our protocols only require logical Clifford gates and Pauli measurements. These operations can be implemented transversally for many stabilizer codes, such as for instance the color codes of Bombin and Martin-Delgado (2006). Thus our protocols are well-suited for networks of small fault-tolerant quantum computers and applications that require fast communication of qubits between $k$ arbitrary pairs of network nodes. A summary of our protocols can be found in Figure 3.
@@ 1.5 Related and follow-up work
Schoute et al. Schoute et al. (2016) consider quantum networks where parties can create EPR-pairs with their immediate neighbors and then use entanglement-swapping combined with efficient routing algorithms to create desired long-distance entanglement. This differs from our approach in allowing the ability to create new EPR-pairs when needed (which requires quantum communication), while we allow only LOCC starting from one fixed entangled resource state.
Hahn, Pappa, and Eisert Hahn et al. (2019) also study a quite similar problem to ours, but starting from a network where some parties are linked via a quantum channel, while some other parties are not (directly) linked at all. In addition to efficiently generating EPR-pairs they also study generating GHZ-states between specified parties.
Pant et al. Pant et al. (2019) study how a network whose nodes are connected via lossy optical links and have limited quantum processing capabilities, can obtain EPR-pairs simultaneously between many pairs of nodes; their limitations per node are analogous to our goal of having only few qubits per party, but they allow quantum communication while we allow only classical communication.
@@
Miguel-Ramiro, Pirker, and Dür Miguel-Ramiro et al. (2023) consider resource states interpolating between the two extreme cases discussed in our introduction: the GHZ state shared among $n$ parties and $\binom{n}{2}$ EPR states shared between each pair of parties. This work proposed clustering and merging algorithms that produce resource states with the desired functionality. However, these methods do not appear to provide $k$-pairable resource states with few qubits per party.
Du, Shang, and Liu Du et al. (2017) study a problem similar to ours but starting from resource states that consist only of pre-shared EPR-pairs between adjacent parties in a given network. Like us, they use entanglement-swapping to create EPR-pairs between distant parties.
Contreras-Tejada, Palazuelos, and de Vicente Contreras-Tejada et al. (2022) gave similar constructions as we gave in Section 2 (with EPR-pairs on the edges of an $n$-vertex graph), but focus primarily on the question for what type of graphs the long-range entanglement survives constant amounts of noise on the edges.
Illiano et al. Illiano et al. (2022) study 1-pairable states with the additional property that the identity of the one pair that ends up sharing an EPR-pair remains unknown to the other $n-2$ parties (in fact one can get this easily from the $n$-party GHZ-state if the other parties broadcast their measurement outcomes to everyone rather than sending it only to the two parties that want an EPR-pair).
Meignant, Markham, and Grosshans Meignant et al. (2019) and Fischer and Townsley Fischer and Townsley (2021) studied what is roughly a partial “dual” of our problem: how many EPR-pairs between which parties of a given $n$-party network are necessary and sufficient to generate a classically given $n$-party graph state?
Dahlberg, Helsen, and Wehner Dahlberg et al. (2020b) show that it is NP-complete to decide whether a classically given $n$-party stabilizer state can be transformed into a set of EPR-pairs on specific qubits using only single-qubit Clifford operations, single-qubit Pauli measurements and classical communication (such protocols are more restricted than the LOCC we allow in our paper). They also give some algorithms to do the transformation in some special cases Dahlberg et al. (2020a).
@@
Our construction in Section 3 has $m=1$ and $k=\Theta(\log n)$, so the number of qubits it uses to achieve $k$-pairability with one qubit per party, scales as $n=2^{\Omega(k)}$. Claudet, Mhalla, and Perdrix Claudet et al. (2023) recently improved this exponential scaling to a polynomial one: they show the existence of $k$-pairable graph states with one qubit per party, using only $n=O(k^{3}(\log k)^{3})$ qubits (=parties). Their proof uses the probabilistic method, so is not as constructive as ours. However, very recently this was improved by Cautrèt, Claudet, Mhalla, Perdrix, Savin, and Thomassé Cautés et al. (2024) to a probabilistic proof with $n=O(k^{2})$ qubits and an explicit construction with $n=O(k^{4})$ qubits.
@@ 2 Constructions with multiple qubits per party
Fact 1 (Broder, Frieze, Upfal Broder et al. (1994)).
@@
It is known Bollobás (1988) that $d$-regular graphs with edge expansion $h>1$ exist for any constant $d\geq 5$ and all large enough $n$. Thus Corollary 2 and Fact 1 imply that for all large enough $n$ there exist $k$-pairable resource states with $k=n/\mbox{\rm polylog}n$ and at most 10 qubits per party.
Let us say that a graph $G$ is path-pairable if the number of vertices $n$ is even and the condition of Lemma 1 holds for $k=n/2$. We shall need the following fact stated as Corollary 2 in Győri et al. (2017).
Fact 2 (Győri, Mezei, Mészáros Győri et al. (2017)).
@@ 3.1 Pairability of CSS stabilizer states
where ${\cal C}\subseteq\mathbb{F}_{2}^{n}$ is a linear code. Such states are known as Calderbank-Shor-Steane (CSS) stabilizer states Calderbank and Shor (1996); Steane (1996); Calderbank et al. (1998). It is well-known that the state $|{\cal C}\rangle$ can be prepared by a quantum circuit of size $O(n^{2})$ for any linear code ${\cal C}$, see for instance Aaronson and Gottesman (2004). We begin by deriving a sufficient condition under which a CSS stabilizer state is $k$-pairable. Below we assume that each of the $n$ parties holds only one qubit.
@@
We assume some familiarity with the stabilizer formalism Calderbank and Shor (1996); Gottesman (1998a); Nielsen and Chuang (2002). Let $\sigma^{x}_{j}$ and $\sigma^{z}_{j}$ be single-qubit Pauli operators acting on the $j$-th qubit tensored with the identity on all other qubits. The resource state $|{\cal C}\rangle$ has Pauli stabilizers [^5] Note that $\sigma^{x}(f)|{\cal C}\rangle=\frac{1}{\sqrt{|{\cal C}|}}\sum_{g\in{\cal C}}\sigma^{x}(f)|g\rangle=\frac{1}{\sqrt{|{\cal C}|}}\sum_{g\in{\cal C}}|f+g\rangle=\frac{1}{\sqrt{|{\cal C}|}}\sum_{g^{\prime}\in{\cal C}}|g^{\prime}\rangle=|{\cal C}\rangle$ for any $f\in{\cal C}$. Here the third equality uses the fact that ${\cal C}$ is a linear code, i.e., a subspace. Furthermore, $\sigma^{z}(\bar{f})|{\cal C}\rangle=|{\cal C}\rangle$ since $\sigma^{z}(\bar{f})|g\rangle=(-1)^{\bar{f}\cdot g}|g\rangle=|g\rangle$ for any $\bar{f}\in{\cal C}^{\perp}$ and $g\in{\cal C}$.
@@
Thus it suffices to show that for any $k\geq 1$ one can choose a sufficiently large $n$ and a linear code ${\cal C}\subseteq\mathbb{F}_{2}^{n}$ that satisfies $k$-pairability conditions CSS1 and CSS2 of Lemma 4. Below we will choose ${\cal C}$ from the family of Reed-Muller codes MacWilliams and Sloane (1977) to achieve this.
@@ 3.2 Reed-Muller codes
For the proof of Facts 3, 4, 5, see e.g. Chapter 13 of MacWilliams and Sloane (1977). As a consequence of Fact 5, the resource state $|{\cal C}\rangle$ with ${\cal C}=\mathsf{RM}(r,m)$ is invariant under a permutation of the $n=2^{m}$ qubits defined as $W_{\varphi}|f\rangle=|f^{\prime}\rangle$, where $f^{\prime}(x)=f(\varphi(x))$ for all $x\in\mathbb{F}_{2}^{m}$. Here $\varphi\,:\,\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{m}$ is any invertible affine map. In other words, $W_{\varphi}|{\cal C}\rangle=|{\cal C}\rangle$. This generalizes the symmetry of the $n$-qubit GHZ-state which is invariant under any permutation of the $n$ qubits.
@@
For the proof see, e.g., Proposition 2 and Corollary 4 in Assmus Jr (1992). We shall see that verification of conditions CSS1 and CSS2 of Lemma 4 with ${\cal C}=\mathsf{RM}(r,m)$ can be reduced to (multiple instances of) the following problem.
@@ 7. I7:
where $H$ is the Hadamard gate. Indeed, it is known Raussendorf et al. (2003) that the graph state $|\psi\rangle$ has stabilizers
@@
For each of the remaining choices of EPR-pairs we numerically examined all $3^{6}$ Pauli measurement bases on qubits $\mathbb{Z}_{10}\setminus\{a_{1},b_{1},a_{2},b_{2}\}$ and computed the final post-measurement state of qubits $a_{1},b_{1},a_{2},b_{2}$ using the standard stabilizer formalism. To test whether the final state is locally equivalent to the desired EPR-pairs, we checked whether the entanglement entropies of the final state obey $S(a_{i})=S(b_{i})=1$ and $S(a_{i}b_{i})=0$ for $i=1,2$. The entanglement entropy of a stabilizer state can be extracted from its tableaux as described in Fattal et al. (2004). Any two-qubit stabilizer state of qubits $a_{i},b_{i}$ satisfying $S(a_{i})=S(b_{i})=1$ and $S(a_{i}b_{i})=0$ has to be maximally entangled and thus equivalent to the EPR-pair modulo single-qubit Clifford gates. We found a Pauli basis generating maximally-entangled states on qubits $\{a_{1},b_{1}\}$ and $\{a_{2},b_{2}\}$ in all considered cases, see Figure 6. We also observed that the graph state $|\psi\rangle$ is not $2$-pairable if the Pauli bases are restricted to $\sigma^{x}$ and $\sigma^{z}$ only.
@@
We verified numerically that no stabilizer state with $n<10$ qubits is $2$-pairable using LOCC protocols based on Pauli measurements, by checking all possible 9-qubit graph states as listed in Adcock et al. (2020). The code is available at https://github.com/yashsharma25/generating-k-epr-pairs
@@ 6 Conclusion and future work
  • Our constructions of $k$-pairable states may be far from optimal in various respects, and it would be interesting to improve them. For example, we already mentioned the follow-up papers Claudet et al. (2023); Cautés et al. (2024) which for the case $m=1$ reduced the exponential number of qubits $n(k)=2^{\Omega(k)}$ of our Reed-Muller-based construction of $k$-pairable states to a polynomial dependence on $k$. The case $m=2$ remains largely open for now, since two qubits per party is not enough to realize entanglement-swapping protocols on expander graphs, see Section 2. All our constructions are based on stabilizer-type resource states. Can one improve the tradeoff between $k$ and $m$ using more general resource states? Can one express the pairability parameter $k$ in terms of some previously studied entanglement measures?
@@
  • How well do our resource states behave under noise? Contreras-Tejada, Palazuelos, and de Vicente Contreras-Tejada et al. (2022) already proved some negative results here for the type of constructions we gave in Section 2 (with EPR-pairs on the edges of an $n$-vertex graph), showing that genuine multipartite entanglement only survives constant amounts of noise per edge if the graph has a lot of connectivity.

### v5

@@ 1.1 Generating EPR-pairs from a resource state
Quantum communication networks combine several quantum computers to enable them to solve interesting tasks from cryptography, communication complexity, distributed computing etc. Building a large-scale quantum communication network is a daunting task that will take many years, but networks with a few small quantum computers are under construction and may start to appear in the next few years [36].
@@
These networks are either based on channels that physically communicate quantum states, or rely on classical communication in tandem with shared entanglement, or a combination of both. Communication over classical channels cannot increase entanglement, so in the absence of quantum channels we have to rely on prior entangled states. For example, if two parties share an EPR-pair, $\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$, then one party can transmit (“teleport”) a qubit to the other via two classical bits of communication, consuming the EPR-pair in the process [4]. If we want to enable many qubits to be sent in this way, then we could start with an entangled state where each pair among the $n$ parties shares its own EPR-pair. This would allow any pair to exchange a qubit, but would require us to start with a rather large initial entangled state of $\binom{n}{2}$ EPR-pairs, and each of the $n$ parties would need to hold $n-1$ qubits (see Figure 1 for $n=4$).
@@ 1.2 Motivation
Multi-party resource states are foundational for tasks such as distributed quantum computing [33, 32, 28], quantum secret sharing [24, 28], and multi-party quantum protocols [12, 19]. One of the most utilized resources in quantum networking is EPR-pairs shared between two parties. Our model aims to capture scenarios involving multiple parties, where it is not known in advance between which parties the quantum resources should be deployed. This situation is similar to a call center, which must dynamically manage incoming calls. Our study concentrates on developing resource states that provide the flexibility to determine the call/connection structure after the resource state has been created.
@@
In this theoretical work we use several simplifying assumptions, most notably that EPR-pairs suffice as the primary resource required for applications. The interesting follow-up work [11, 10], discussed at the end of this section, subsequently removed this assumption.
@@ 1.3 Our results 1: constructions of $k$-pairable resource states
Our protocols for generating EPR-pairs can be made fault-tolerant if each party encodes their $m$ qubits by some stabilizer-type quantum code [21]. Importantly, the encoded versions of our protocols only require logical Clifford gates and Pauli measurements. These operations can be implemented transversally for many stabilizer codes, such as for instance the color codes of [6]. Thus our protocols are well-suited for networks of small fault-tolerant quantum computers and applications that require fast communication of qubits between $k$ arbitrary pairs of network nodes. A summary of our protocols can be found in Figure 3.
@@ 1.5 Related and follow-up work
Schoute et al. [34] consider quantum networks where parties can create EPR-pairs with their immediate neighbors and then use entanglement-swapping combined with efficient routing algorithms to create desired long-distance entanglement. This differs from our approach in allowing the ability to create new EPR-pairs when needed (which requires quantum communication), while we allow only LOCC starting from one fixed entangled resource state.
Hahn, Pappa, and Eisert [23] also study a quite similar problem to ours, but starting from a network where some parties are linked via a quantum channel, while some other parties are not (directly) linked at all. In addition to efficiently generating EPR-pairs they also study generating GHZ-states between specified parties.
Pant et al. [31] study how a network whose nodes are connected via lossy optical links and have limited quantum processing capabilities, can obtain EPR-pairs simultaneously between many pairs of nodes; their limitations per node are analogous to our goal of having only few qubits per party, but they allow quantum communication while we allow only classical communication.
@@
Miguel-Ramiro, Pirker, and Dür [29] consider resource states interpolating between the two extreme cases discussed in our introduction: the GHZ state shared among $n$ parties and $\binom{n}{2}$ EPR states shared between each pair of parties. This work proposed clustering and merging algorithms that produce resource states with the desired functionality. However, these methods do not appear to provide $k$-pairable resource states with few qubits per party.
Du, Shang, and Liu [16] study a problem similar to ours but starting from resource states that consist only of pre-shared EPR-pairs between adjacent parties in a given network. Like us, they use entanglement-swapping to create EPR-pairs between distant parties.
Contreras-Tejada, Palazuelos, and de Vicente [13] gave similar constructions as we gave in Section 2 (with EPR-pairs on the edges of an $n$-vertex graph), but focus primarily on the question for what type of graphs the long-range entanglement survives constant amounts of noise on the edges.
Illiano et al. [25] study 1-pairable states with the additional property that the identity of the one pair that ends up sharing an EPR-pair remains unknown to the other $n-2$ parties (in fact one can get this easily from the $n$-party GHZ-state if the other parties broadcast their measurement outcomes to everyone rather than sending it only to the two parties that want an EPR-pair).
Meignant, Markham, and Grosshans [27] and Fischer and Townsley [18] studied what is roughly a partial “dual” of our problem: how many EPR-pairs between which parties of a given $n$-party network are necessary and sufficient to generate a classically given $n$-party graph state?
Dahlberg, Helsen, and Wehner [15] show that it is NP-complete to decide whether a classically given $n$-party stabilizer state can be transformed into a set of EPR-pairs on specific qubits using only single-qubit Clifford operations, single-qubit Pauli measurements and classical communication (such protocols are more restricted than the LOCC we allow in our paper). They also give some algorithms to do the transformation in some special cases [14].
@@
Our construction in Section 3 has $m=1$ and $k=\Theta(\log n)$, so the number of qubits it uses to achieve $k$-pairability with one qubit per party, scales as $n=2^{\Omega(k)}$. Claudet, Mhalla, and Perdrix [11] recently improved this exponential scaling to a polynomial one: they show the existence of $k$-pairable graph states with one qubit per party, using only $n=O(k^{3}(\log k)^{3})$ qubits (=parties). Their proof uses the probabilistic method, so is not as constructive as ours. However, very recently this was improved by Cautrèt, Claudet, Mhalla, Perdrix, Savin, and Thomassé [10] to a probabilistic proof with $n=O(k^{2})$ qubits and an explicit construction with $n=O(k^{4})$ qubits.
@@ 2 Constructions with multiple qubits per party
Fact 1 (Broder, Frieze, Upfal [7]).
@@
It is known [5] that $d$-regular graphs with edge expansion $h>1$ exist for any constant $d\geq 5$ and all large enough $n$. Thus Corollary 2 and Fact 1 imply that for all large enough $n$ there exist $k$-pairable resource states with $k=n/\mbox{\rm polylog}n$ and at most 10 qubits per party.
Let us say that a graph $G$ is path-pairable if the number of vertices $n$ is even and the condition of Lemma 1 holds for $k=n/2$. We shall need the following fact stated as Corollary 2 in [22].
Fact 2 (Győri, Mezei, Mészáros [22]).
@@ 3.1 Pairability of CSS stabilizer states
where ${\cal C}\subseteq\mathbb{F}_{2}^{n}$ is a linear code. Such states are known as Calderbank-Shor-Steane (CSS) stabilizer states [9, 35, 8]. It is well-known that the state $|{\cal C}\rangle$ can be prepared by a quantum circuit of size $O(n^{2})$ for any linear code ${\cal C}$, see for instance [1]. We begin by deriving a sufficient condition under which a CSS stabilizer state is $k$-pairable. Below we assume that each of the $n$ parties holds only one qubit.
@@
We assume some familiarity with the stabilizer formalism [9, 20, 30]. Let $\sigma^{x}_{j}$ and $\sigma^{z}_{j}$ be single-qubit Pauli operators acting on the $j$-th qubit tensored with the identity on all other qubits. The resource state $|{\cal C}\rangle$ has Pauli stabilizers [^5] Note that $\sigma^{x}(f)|{\cal C}\rangle=\frac{1}{\sqrt{|{\cal C}|}}\sum_{g\in{\cal C}}\sigma^{x}(f)|g\rangle=\frac{1}{\sqrt{|{\cal C}|}}\sum_{g\in{\cal C}}|f+g\rangle=\frac{1}{\sqrt{|{\cal C}|}}\sum_{g^{\prime}\in{\cal C}}|g^{\prime}\rangle=|{\cal C}\rangle$ for any $f\in{\cal C}$. Here the third equality uses the fact that ${\cal C}$ is a linear code, i.e., a subspace. Furthermore, $\sigma^{z}(\bar{f})|{\cal C}\rangle=|{\cal C}\rangle$ since $\sigma^{z}(\bar{f})|g\rangle=(-1)^{\bar{f}\cdot g}|g\rangle=|g\rangle$ for any $\bar{f}\in{\cal C}^{\perp}$ and $g\in{\cal C}$.
@@
Thus it suffices to show that for any $k\geq 1$ one can choose a sufficiently large $n$ and a linear code ${\cal C}\subseteq\mathbb{F}_{2}^{n}$ that satisfies $k$-pairability conditions CSS1 and CSS2 of Lemma 4. Below we will choose ${\cal C}$ from the family of Reed-Muller codes [26] to achieve this.
@@ 3.2 Reed-Muller codes
For the proof of Facts 3, 4, 5, see e.g. Chapter 13 of [26]. As a consequence of Fact 5, the resource state $|{\cal C}\rangle$ with ${\cal C}=\mathsf{RM}(r,m)$ is invariant under a permutation of the $n=2^{m}$ qubits defined as $W_{\varphi}|f\rangle=|f^{\prime}\rangle$, where $f^{\prime}(x)=f(\varphi(x))$ for all $x\in\mathbb{F}_{2}^{m}$. Here $\varphi\,:\,\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{m}$ is any invertible affine map. In other words, $W_{\varphi}|{\cal C}\rangle=|{\cal C}\rangle$. This generalizes the symmetry of the $n$-qubit GHZ-state which is invariant under any permutation of the $n$ qubits.
@@
For the proof see, e.g., Proposition 2 and Corollary 4 in [3]. We shall see that verification of conditions CSS1 and CSS2 of Lemma 4 with ${\cal C}=\mathsf{RM}(r,m)$ can be reduced to (multiple instances of) the following problem.
@@ 7. I7:
where $H$ is the Hadamard gate. Indeed, it is known [33] that the graph state $|\psi\rangle$ has stabilizers
@@
For each of the remaining choices of EPR-pairs we numerically examined all $3^{6}$ Pauli measurement bases on qubits $\mathbb{Z}_{10}\setminus\{a_{1},b_{1},a_{2},b_{2}\}$ and computed the final post-measurement state of qubits $a_{1},b_{1},a_{2},b_{2}$ using the standard stabilizer formalism. To test whether the final state is locally equivalent to the desired EPR-pairs, we checked whether the entanglement entropies of the final state obey $S(a_{i})=S(b_{i})=1$ and $S(a_{i}b_{i})=0$ for $i=1,2$. The entanglement entropy of a stabilizer state can be extracted from its tableaux as described in [17]. Any two-qubit stabilizer state of qubits $a_{i},b_{i}$ satisfying $S(a_{i})=S(b_{i})=1$ and $S(a_{i}b_{i})=0$ has to be maximally entangled and thus equivalent to the EPR-pair modulo single-qubit Clifford gates. We found a Pauli basis generating maximally-entangled states on qubits $\{a_{1},b_{1}\}$ and $\{a_{2},b_{2}\}$ in all considered cases, see Figure 6. We also observed that the graph state $|\psi\rangle$ is not $2$-pairable if the Pauli bases are restricted to $\sigma^{x}$ and $\sigma^{z}$ only.
@@
We verified numerically that no stabilizer state with $n<10$ qubits is $2$-pairable using LOCC protocols based on Pauli measurements, by checking all possible 9-qubit graph states as listed in [2]. The code is available at https://github.com/yashsharma25/generating-k-epr-pairs
@@ 6 Conclusion and future work
  • Our constructions of $k$-pairable states may be far from optimal in various respects, and it would be interesting to improve them. For example, we already mentioned the follow-up papers [11, 10] which for the case $m=1$ reduced the exponential number of qubits $n(k)=2^{\Omega(k)}$ of our Reed-Muller-based construction of $k$-pairable states to a polynomial dependence on $k$. The case $m=2$ remains largely open for now, since two qubits per party is not enough to realize entanglement-swapping protocols on expander graphs, see Section 2. All our constructions are based on stabilizer-type resource states. Can one improve the tradeoff between $k$ and $m$ using more general resource states? Can one express the pairability parameter $k$ in terms of some previously studied entanglement measures?
@@
  • How well do our resource states behave under noise? Contreras-Tejada, Palazuelos, and de Vicente [13] already proved some negative results here for the type of constructions we gave in Section 2 (with EPR-pairs on the edges of an $n$-vertex graph), showing that genuine multipartite entanglement only survives constant amounts of noise per edge if the graph has a lot of connectivity.